Question 1) Mack’s guitar fabrication shop produces low-cost, highly durable guitars for beginners. Typically, out of the 100 guitars that begin production each month, only 80 percent are considered good enough to sell. The other 20 percent are scrapped due to quality problems that are identified after they have completed the production process. Each guitar sells for $250. Because some of the production process is automated, each guitar only requires 10 labor hours. Each employee works an average 160 hours per month. Labor is paid at $10/hour, materials cost is $40/guitar, and overhead is $4,000.

  1. Calculate the labor and multifactor productivity ratios.
  2. After some study, the operations manager Darren Funk recommends three options to improve the company’s multifactor productivity: (1) increase the sales price by 10 percent, (2) improve quality so that only 10 percent are defective, or (3) reduce labor, material, and overhead costs by 10 percent. Which option has the greatest impact on the multifactor productivity measure?
Answer #1

a)

Output = 100 guitars produced * 80% good

= 80 guitars

Labor input = 100 guitars * 10 hours per guitar

= 1000 hours

Labor productivity ratio = Output / Labor input

= 80 / 1000

= 0.08 guitars per labor hour

----------------------------------------------------

Dollar output = 80 guitars * $ 250 per guitar

= $ 20,000

Total input cost = Labor cost + Materials cost + Overhead cost

= 1000*$10 + 100*$40 + $4000

= $ 18,000

Multifactor productivity ratio = Output / Total input cost

= 20000 / 18000

= 1.111

===============================

b)

(1) Increase sales price by 10 percent

New sales price = 250*1.1 = $ 275

New output = 80*275 = $ 22,000

New multifactor productivity ratio = 22000/18000 = 1.222

(2) Only 10 percent defective

Good guitars produced = 100*90% = 90

New output = 90*250 = $ 22,500

New multifactor productivity ratio = 22500/18000 = 1.25

(3) Reduce labor, material and overhead cost by 10 percent

New total cost = 18000*(1-10%) = $ 16,200

New multifactor productivity ratio = 20000/16200 = 1.235

Conclusion: We see that mltifactor productivity ratio is the highest for option (2) , i.e. improve quality so that only 10 percent are defective